# Analyzing Variations in Dependency Distributions Due to Code Smell Interactions

### v3

@@ 4. Results
This section presents our results. Due to space constraints, we include only the most interesting results. Our replication package provides the complete dataset (blind_2025_15624918, 17).
@@ 4.1. Interaction Frequency
RQ1 Summary: Overall, most code smell pairs frequently interact with one another, with over 50% of the pairs exhibiting an interaction frequency above 40%. This highlights the importance of evaluating the consequences of code smell interactions.
@@ 4.2. General Analysis Union Dataset
Figure 3 presents our general analysis results for the Union dataset. The y-axis shows the 10 code smell pairs, while the x-axis shows the four contrasts. Each cell contains three pieces of information: Cliff’s delta, its interpretation, and the median value for the two interaction types being compared (in parentheses). The gradient scale of the cell color is based on Cliff’s delta, ranging from red (1) to grey (0) to blue (-1). Cells in bold indicate a significant Mann-Whitney U test, where we can reject $H_{0}$.
@@
Finding 1  Code smell interactions are associated with a small to medium, statistically significant increase in the total number of dependencies in 28 out of 36 possible cases.
@@ 4.3. Specific Analysis Union Dataset
(use, FM, Field) 10 28
(call, FM, Constructor) 6 33
(create, FM, Class) 6 33
(parameter, FM, Class) 4 24
(return, FM, Class) 4 26
@@
Finding 2  Code smell interactions are frequently associated with significant differences in data access and FM call dependencies, observed in 25%–67% of tests with at least a small effect size, and up to 93% when the small-effect-size threshold is removed.
@@
Finding 3  For a given relative location and role, most code smells show a consistent change direction in the number of data access and FM call dependencies when interacting with another code smell, although the direction varies across dependency types, locations, and roles.
4.4. General and Specific Analysis Per System
@@
Figure 8 presents two heatmaps for the per-system general analysis. We include only the projects for which the statistical test can be conducted, as discussed in Section 3.6. The left heatmap shows the significance rate, defined as the percentage of projects that yield a statistically significant result. The right heatmap shows the consistency score, calculated as the difference between the number of projects with a nonnegative Cliff’s delta and those with a negative Cliff’s delta, divided by the total number of included projects. The score ranges from -1 to 1 and indicates whether projects tend to share the same change direction. A value near 0 indicates low consistency, whereas values near -1 or 1 indicate perfect consistency (all increases or all decreases, respectively). The significance rate ranges from 0% to 82%, and the consistency score ranges from -0.44 to 1.00, indicating that, for some code smell pairs, different projects exhibit mixed significance levels and opposite change directions.
@@
Consider the cell at row (GC, DC) and column (nonCS2, Diff. Class) in Figure 8, which compares GC-DC with GC-nonDC. It has a significance rate of $49\%$ and a consistency score of $0.71$. This indicates that about half of the projects are not significant, with a tendency toward a nonnegative Cliff’s delta. Figure 9 presents the corresponding per system specific analysis for this pair, focusing on FM call and data access dependencies when the artifacts are in different classes. For all dependency types in both flow directions, over $75\%$ of individual projects follow the same change direction observed in the Union dataset (Figures 4-7). Most projects that do not follow this pattern are either outliers, yield non-significant results, or contain limited data that reduce the power of statistical tests. We also observe that Cliff’s delta for the forward (call, FM, Accessor) dependency tends to be positive, whereas the other three tend to be negative, with the change magnitude varying across projects. These opposite change directions among dependency types, along with the varying magnitudes across projects, help explain why, in the per system general analysis, certain contrasts exhibit low significance rates and consistency across projects.
Finding 4  Across projects, change magnitudes vary for both data access and FM call dependencies, yet the change direction remains largely consistent and aligns with the trend observed in the Union dataset.
RQ2 Summary: Overall, code smell interactions are generally associated with an increase in the total number of dependencies. Significant changes with at least a small effect size are most frequently observed in data access and FM call dependencies. For a given relative location and role, individual code smells tend to exhibit a consistent change direction in these dependency types when interacting with other smells.
@@ 5.2. Design Insights Derived from Dependency Distribution
FE and BM are weak FM providers, meaning their functionality is less likely to be relied upon during interactions compared to non-code smell artifacts. The recommendation is to create small methods, not because large ones are harder to understand, but because as methods grow, they often take on multiple responsibilities that are not needed simultaneously, reducing their usefulness as single-responsibility units for other artifacts. In contrast, GC and BC are strong FM providers and consumers, invoking and offering many FMs due to their large and multi-responsibility nature. Therefore, creating small, single-responsibility classes is advisable.
@@ 5.3. Interaction-Based Inference of Code Smells
The results suggest that code smells frequently interact with each other, especially those involving GC and DC (Table 3). Therefore, examining artifacts that interact with confirmed code smells may help identify additional code smells more confidently and efficiently. We propose a potential approach to achieve this, based on the typology. The core idea is that when the existence of certain code smells (e.g., manually confirmed ones) is known, but others remain uncertain, the interactions between them can provide useful evidence. Specifically, if the expected change direction from the typology aligns with the observed change direction in the number of dependencies, we can increase our confidence in the presence of the uncertain code smell.
For example, consider a GC that has been manually confirmed. We examine all artifacts interacting with this GC and identify a possible FE. To increase our confidence, we can count the number of FM calls for all methods interacting with the GC. If the number of FM calls from the GC to the suspected FE is relatively low, this provides supporting evidence. Since GC is a strong FM consumer, we would typically expect a high number of FM call dependencies when it interacts with other artifacts. However, the FE is a weak FM provider, offering fewer FMs compared to nonFE artifacts that are likely present in other methods interacting with the GC. Thus, a low number of FM calls between the GC and the suspected FE supports the existence of the FE.
Furthermore, interactions involving additional confirmed code smells can strengthen our confidence. For instance, if a confirmed DC also interacts with the FE, a relatively high number of data access dependencies from the FE to the DC, compared with those from the FE to other classes, would support our speculation. Since FE is a strong data consumer and DC is a strong data provider, such interactions are expected to exhibit higher data access dependency.
@@ 5.4. Interaction-Guided Refactoring Strategies
For example, when FE interacts with DC, we expect a higher than usual number of data access dependencies from FE to DC. This suggests a potential refactoring strategy: move FE into DC. Doing so would prevent FE from envying DC’s data and would add a functional method to DC, making it less of a pure data holder. In Grasscutter [^4] A game server: https://github.com/Grasscutters/Grasscutter, Friendship#toProto (an FE) heavily accesses the data of PlayerProfile (a DC) to serialize it, suggesting that Friendship#toProto could be moved into PlayerProfile to reduce foreign data access and make the DC more behaviorally rich.
Another example involves GC interacting with BC, where we observe a higher than usual number of data access and FM call dependencies. Both BC and GC represent large, complex classes that often contain multiple responsibilities. A common refactoring strategy is to extract classes from them, each representing a single responsibility (lacerda2020code, 58). However, the extensive dependencies between them suggest that the responsibilities may be interwoven across the two classes. Therefore, instead of extracting separate classes from each, it may be better to extract a shared class from both, considering their interactions. In Grasscutter, methods such as upgradeWeapon, promoteWeapon, and calcWeaponUpgradeReturnItems in InventorySystem (a BC) repeatedly operate on the same GameItem’s (a GC) fields, such as level and exp, and always conclude with a call to save. Extracting these methods and the related data into a dedicated WeaponProgressionService would create a cohesive unit focused on weapon advancement, reducing foreign data access and coupling, and decreasing the size and complexity of both original classes.
@@ 5.5. Threats to Validity
The choice of detection tool represents a potential threat to construct validity. Many existing tools lack clear documentation of their detection heuristics and code smell definitions, and they often support only a limited subset of smells. To address this, we developed our own tool based on the metric-based detection framework proposed by Lanza and Marinescu (lanza2007object, 13) and made it publicly available to enhance transparency and replicability.
Another threat to internal validity arises from detection errors. We mitigated this by manually verifying metric measurements and the detected code smell instances, which showed a low false-positive rate. However, we did not evaluate false negatives due to the large number of non-smelly artifacts and the higher cognitive effort required for their inspection. While our goal was to ensure that analyzed smells were genuine and that observed trends were reliable, assessing false negatives remains an important task for future work.
A threat to internal validity in our study arises from how we define non-code smell artifacts. Specifically, a nonCS1 artifact may still exhibit other code smells such as CS2, which could introduce confounding effects when comparing interaction cases. However, as shown in Table 2, smelly artifacts constitute only a small fraction of the total dataset. Consequently, a large portion of artifacts are clean and included in the non-code smell category for any smell type. Therefore, the potential bias introduced by this overlap is expected to be minimal, and our findings remain robust.
6. Conclusions
@@
Empirical evidence indicates that most pairs of code smells tend to interact frequently, and these interactions are associated with a significant increase in the total number of dependencies. Although this outcome may not be unexpected, to the best of our knowledge, this study provides the first rigorous empirical evidence to substantiate this relationship. Moreover, code smell interactions exhibit consistent change direction in certain dependency types. These findings offer meaningful design insights, suggesting that practitioners should prioritize resolving interacting code smells rather than those occurring in isolation. Interacting smells not only display denser and more complex dependency structures but also show consistent patterns of dependency change that may reflect deeper design issues. At the same time, this information can guide more effective refactoring strategies and inform the development of improved code smell detection techniques.
Future work will investigate a broader range of code smells, interactions among code smells of the same type, and the influence of these interactions on other maintainability metrics. Additionally, future studies could examine contextual factors, such as project domain and programming language, to assess their impact on code smell interactions.

### v4

@@ 4. Results
This section presents our results. Due to space constraints, we include only the most interesting results. Our replication package provides the complete dataset (Zhang et al., 2025).
@@ 4.1. Interaction Frequency
[…]
@@ 4.2. General Analysis Union Dataset
Figure 3 presents our general analysis results for the Union dataset (a version of this figure appears in our Doctoral Symposium paper (Zhang, 2025)). The y-axis shows the 10 code smell pairs, while the x-axis shows the four contrasts. Each cell contains three pieces of information: Cliff’s delta, its interpretation, and the median value for the two interaction types being compared (in parentheses). The gradient scale of the cell color is based on Cliff’s delta, ranging from red (1) to grey (0) to blue (-1). Cells in bold indicate a significant Mann-Whitney U test, where we can reject $H_{0}$.
@@
Cliff’s Delta Heatmap (bold = significant Mann-Whitney U test)
@@
[…]
Sensitivity Analysis. By definition, a nonCS1 artifact may still exhibit other code smells such as CS2, which could introduce confounding effects and compromise the interpretability of the results. We conducted a sensitivity analysis by excluding nonCS artifacts that exhibit any of the other four smells (e.g., a nonCS1 artifact is already free of its own smell type, CS1, so only the other four need to be checked). The output confirms that the percentage of nonCS artifacts that are not completely clean is very low: 5.56% for nonBC, 0.31% for nonBM, 2.14% for nonDC, 0.27% for nonFE, and 4.37% for nonGC. We reran the general analysis on the union dataset using this clean subset and regenerated Figure 3. The results indicate that the effect sizes for most cells increased slightly, and a few contrasts that were previously not significant became significant. Therefore, excluding nonCS instances that are not completely clean does not invalidate our claims; rather, it further strengthens them.
@@ 4.3. Specific Analysis Union Dataset
(use, FM, Field) 10 27
(call, FM, Constructor) 6 32
(create, FM, Class) 6 32
(parameter, FM, Class) 4 23
(return, FM, Class) 4 25
@@
[…]
@@
[…]
4.4. Per-System General and Specific Analyses
@@
Significance rate and consistency score across all projects
Figure 8 presents two heatmaps for the per-system general analysis. We include only the projects for which the statistical test can be conducted, as discussed in Section 3.6. The left heatmap shows the significance rate, defined as the percentage of projects that yield a statistically significant result. The right heatmap shows the consistency score, calculated as the difference between the number of projects with a nonnegative Cliff’s delta and those with a negative Cliff’s delta, divided by the total number of included projects. The score ranges from -1 to 1 and indicates whether projects tend to share the same change direction. A value near 0 indicates low consistency, whereas values near -1 or 1 indicate perfect consistency (all increases or all decreases, respectively). The significance rate ranges from 0% to 76%, and the consistency score ranges from -0.44 to 1.00, indicating that, for some code smell pairs, different projects exhibit mixed significance levels and opposite change directions.
@@
[…]
Consider the cell at row (GC, DC) and column (nonCS2, Diff.) in Figure 8, which compares GC-DC with GC-nonDC. It has a significance rate of $43\%$ and a consistency score of $0.71$. This indicates that about half of the projects are not significant, with a tendency toward a nonnegative Cliff’s delta. Figure 9 presents the corresponding per-system specific analysis for this pair, focusing on FM call and data access dependencies when the artifacts are in different classes. For all dependency types in both flow directions, over $75\%$ of individual projects follow the same change direction observed in the Union dataset (Figures 4-7). Most projects that do not follow this pattern are either outliers, yield non-significant results, or contain limited data that reduce the power of statistical tests. We also observe that Cliff’s delta for the forward (call, FM, Accessor) dependency tends to be positive, whereas the other three tend to be negative, with the change magnitude varying across projects. These opposite change directions among dependency types, along with the varying magnitudes across projects, help explain why, in the per-system general analysis, certain contrasts exhibit low significance rates and consistency across projects.
[…]
@@ 5.1. A Multi-Dimensional Typology
A Multi-Dimensional Typology Based on Dependency Type and Role
@@ 5.2. Design Insights Derived from Dependency Distribution
FE and BM are weak FM providers, meaning their functionality is less likely to be relied upon during interactions compared to non-code smell artifacts. Therefore, it is recommended to create small methods, not only because large ones are harder to understand, but also because as methods grow, they often take on multiple responsibilities that are not needed simultaneously, reducing their usefulness as single-responsibility units for other artifacts. In contrast, GC and BC are strong FM providers and consumers, invoking and offering many FMs due to their large and multi-responsibility nature. Therefore, creating small, single-responsibility classes is advisable.
@@ 5.3. Interaction-Based Inference of Code Smells
It has been suggested that the presence of certain types of smells can support the existence of other types of smells (Pietrzak and Walter, 2006) (e.g., the existence of data classes supports the existence of feature envy methods that use their data). Such information can improve the efficiency and effectiveness of detecting additional code smells when certain code smells have been confirmed (e.g., manually confirmed). We can also leverage interaction information and the typology to offer another perspective on evidence of support.
Given a confirmed code smell in one artifact, we assess whether a related artifact exhibits a different smell by examining their interactions. Specifically, if the expected change direction derived from the typology aligns with the observed change direction in the number of dependencies, we can increase our confidence that the related artifact exhibits the code smell.
For example, consider a GC that has been manually confirmed. Suppose there is a method $m_{1}$ for which we want to detect whether it exhibits FE. If $m_{1}$ does not interact with the GC, then we cannot utilize interaction information. However, if they do interact, one useful interaction signal is the number of FM call dependencies from the GC to each method it interacts with, including $m_{1}$. If the number of FM calls from the GC to $m_{1}$ is relatively low compared to those to other methods that the GC interacts with, this provides supporting evidence that $m_{1}$ exhibits FE. Since a GC is a strong FM consumer, we would typically expect a high number of FM calls when it interacts with other artifacts. However, if $m_{1}$ exhibits FE, then it is a weak FM provider and will be called less often (i.e., it will receive fewer FM calls) compared to nonFE methods that the GC interacts with. Thus, a low number of FM calls from the GC to $m_{1}$ supports the existence of FE in $m_{1}$.
Furthermore, interactions involving additional confirmed code smells can further strengthen our confidence. For instance, if a confirmed DC also interacts with $m_{1}$, a relatively high number of data access dependencies from $m_{1}$ to the DC, compared with those from other methods that interact with the DC, would further support that $m_{1}$ exhibits FE. If $m_{1}$ exhibits FE, then it is a strong data consumer and the DC is a strong data provider; therefore, such interactions are expected to exhibit a high number of data access dependencies.
@@ 5.4. Interaction-Guided Refactoring Strategies
For example, when FE interacts with DC, we expect a higher-than-usual number of data access dependencies from FE to DC. This suggests a potential refactoring strategy: move FE into DC. Doing so would prevent FE from envying DC’s data and would add a functional method to DC, making it less of a pure data holder. In Grasscutter [^4] Grasscutter is an open-source anime game server emulator. Source code: https://github.com/Grasscutters/Grasscutter, Friendship#toProto (an FE) heavily accesses the data of PlayerProfile (a DC) to serialize it, suggesting that Friendship#toProto could be moved into PlayerProfile to reduce foreign data access and make the DC more behaviorally rich.
Another example involves GC interacting with BC, where we observe a higher-than-usual number of data access and FM call dependencies. Both BC and GC represent large, complex classes that often contain multiple responsibilities. A common refactoring strategy is to extract classes from them, each representing a single responsibility (Lacerda et al., 2020). However, the extensive dependencies between them suggest that the responsibilities may be interwoven across the two classes. Therefore, instead of extracting separate classes from each, it may be better to extract a shared class from both, considering their interactions. In Grasscutter, methods such as upgradeWeapon, promoteWeapon, and calcWeaponUpgradeReturnItems in InventorySystem (a BC) repeatedly operate on the same GameItem’s (a GC) fields, such as level and exp, and always conclude with a call to save. Extracting these methods and the related data into a dedicated WeaponProgressionService would create a cohesive unit focused on weapon advancement, reducing foreign data access and coupling, and decreasing the size and complexity of both original classes.
@@ 5.5. Threats to Validity
The choice of detection tool represents a potential threat to construct validity. Many existing tools lack clear documentation of their detection heuristics and code smell definitions, and they often support only a limited subset of smells. To address this, we developed our own tool based on the metric-based detection framework proposed by Lanza and Marinescu (Lanza and Marinescu, 2007) and made it publicly available to enhance transparency and replicability.
A threat to internal validity arises from our use of stricter detection thresholds to minimize false positive rates; as a consequence, false negative rates may increase. To assess this threat, we conducted a manual evaluation, which indicates that both false positive and false negative rates are low, although false negative rates are slightly higher. Undetected smell instances would be misclassified into non-smelly groups, diluting the contrast between interaction types and biasing effect sizes toward zero. However, given the overall low false positive and false negative rates, we expect this threat to be minimal.
Another threat to internal validity arises from how we define non-code smell artifacts: a nonCS1 artifact may still exhibit other smells (e.g., CS2), which could potentially introduce confounding effects. To assess this threat, we conducted a sensitivity analysis by excluding nonCS artifacts that exhibit any of the other four smells and reran the analysis. The results show that such contamination is rare and that our main findings remain consistent.
A further threat to internal validity is that inherently different dependency levels between code smell and non-code smell artifacts may confound our comparisons; that is, the observed differences in dependency distributions may be caused by differences in artifact dependency levels rather than the interaction itself. Since our study is observational, we do not claim causal effects. We interpret the results as associations between smell interactions and dependency distributions.
6. Conclusions and Future Work
@@
Our results suggest that most pairs of code smells interact frequently, and that these interactions are associated with a significant increase in the total number of dependencies. Although this outcome may not be unexpected, to the best of our knowledge, this study provides the first rigorous empirical evidence to substantiate this relationship. Moreover, code smell interactions exhibit consistent change direction in certain dependency types. These findings offer meaningful design insights, suggesting that practitioners should prioritize resolving interacting code smells rather than those occurring in isolation. Interacting smells not only display denser and more complex dependency structures but also show consistent patterns of dependency change that may reflect deeper design issues. At the same time, this information can guide more effective refactoring strategies and inform the development of improved code smell detection techniques.
Future work could investigate a broader range of code smells, interactions among code smells of the same type, and the influence of these interactions on other maintainability metrics. Additionally, contextual factors such as project domain and programming language could be examined to assess their impact on code smell interactions. Furthermore, interaction effects could be better isolated to determine whether smell interactions have causal impacts on dependency distributions.
Future work could also formulate and empirically test hypotheses that link dependency distribution patterns to process outcomes (e.g., change proneness, churn, or effort). While this study does not directly evaluate process outcomes, our results on dependency distribution patterns provide a foundational characterization that can inform such analyses. For example, increased dependencies between interacting smelly artifacts may create additional pathways for change propagation and thus be associated with higher change proneness and maintenance effort.